# SchurVINS: Schur Complement-Based Lightweight Visual Inertial Navigation System

### v4

@@ 3.1 State Definition
where {G}, {I} and {C} are the global frame, local frame and camera frame, respectively. ${{}^{G}\textbf{p}_{I}}$ and ${{}^{G}\textbf{v}_{I}}$ are the position and velocity of IMU expressed in {G}, respectively. ${{}^{G}_{I}\textbf{q}}$ represents the rotation quaternion from {I} to {G} (in this paper, quaternion obeys Hamilton rules [28]). The vectors $\textbf{b}_{a}$ and $\textbf{b}_{g}$ individually represent the biases of the angular velocity and linear acceleration from the IMU. And thus the corresponding EKF error-state of ${\textbf{x}_{I}}$ is defined as Eq. 2
@@ 3.2 Propagation and Augmentation
$\begin{matrix}{}^{G}_{I}\dot{\hat{\textbf{q}}}=\frac{1}{2}{{}^{G}_{I}{\hat{\textbf{q}}}}\otimes\Omega(\hat{\boldsymbol{\omega}})\end{matrix}$ (5)
@@
where $\hat{\boldsymbol{\omega}}=\boldsymbol{\omega}_{m}-\hat{\textbf{b}}_{g}$ and $\hat{\textbf{a}}=\textbf{a}_{m}-\hat{\textbf{b}}_{a}$ are IMU measurements with biases discarded, and
$\Omega{(\hat{\boldsymbol{\omega}})}=\begin{pmatrix}0&-{\hat{\boldsymbol{\omega}}}^{\mathsf{T}}\\ \hat{\boldsymbol{\omega}}&-[\hat{\boldsymbol{\omega}}]_{\times}\end{pmatrix}$
where $[\hat{\boldsymbol{\omega}}]_{\times}$ is skew symmetric matrix of $\hat{\boldsymbol{\omega}}$.
Based on Eqs.(5) to (7), the linearized continuous dynamics for the error IMU state is defined as
@@ 3.3 Schur Complement-Based State Update
$\begin{matrix}\textbf{r}=\begin{bmatrix}\textbf{J}_{x}&\textbf{J}_{f}\end{bmatrix}\begin{bmatrix}{\tilde{\textbf{X}}}\\ {{}^{G}{\tilde{\textbf{p}}_{f}}}\end{bmatrix}+\textbf{n},&\textbf{n}=\begin{bmatrix}u,u,u,\dotsb,u\end{bmatrix}^{\mathsf{T}}\end{matrix}$ (19)
where, r and $\begin{bmatrix}\textbf{J}_{x}&\textbf{J}_{f}\end{bmatrix}$ are respectively the stacked residual and stacked Jacobian. $\textbf{J}_{x}$ and $\textbf{J}_{f}$ are jacobian with respect to states and landmark positions, respectively. n is the stacked measurement noise, and the measurement covariance of n is $\textbf{R}=diag(u^{2},u^{2},\dotsb,u^{2})$, where $u$ is one of standard deviation of n.
Unlike [29, 10, 7], in this paper, the residual model Eq. 19 is projected into the jacobian space $\begin{bmatrix}\textbf{J}_{x}&\textbf{J}_{f}\end{bmatrix}^{\mathsf{T}}$for formulating equivalent residual equations, which consist of gradient and hessian and observation covariance shown in Eqs. (20) and (21) below. It is worth highlighting that this strategy is an alternative to QR decomposition strategy [22] for speeding-up in any EKF systems with hyper high-dimensional measurements.
@@ 3.4 EKF-based Landmark Solver
Furthermore, it is an alternative method that the resulting $\tilde{\textbf{X}}$ could also be back substituted into the general residual model  Eq. 16 to establish the alternative equivalent residual equation for landmark.
@@ 4.1 Accuracy
In Tab. 1, our method obtains the lowest average RMSE in filter-based methods reported on the dataset so far, as well as outperforms the majority of optimization-based methods. Besides, our approach obtains the similar accuracy with wellknown optimization-based method BASALT and slightly lower accuracy than the recent competitor DM-VIO. The re-evaluation experiment in Tab. 2 is within expectation absolutely. It is worth highlighting that, although degrading in accuracy slightly compared with the two optimization-based competitors, our method achieves obviously lower computational complexity than both of them with details in the next subsection.
@@ 4.2 Efficiency
As demonstrated in Tab. 3, SchurVINS achieves almost the lowest processor usage compared with all the mentioned VINS algorithms. Especially, SVO2.0-wo requires similar cpu usage with SchurVINS, but it suffers from notable inaccuracy since it is almost pure Visual Odometry (VO). To thoroughly investigate the underlying reasons contributing to the efficiency advantages of SchurVINS, the experiment to meticulously analyze the overhead time of SchurVINS including the comparison to both SVO2.0 and the widely-recognized filter-based OpenVINS is carried out in Tab. 4.

### v5

@@ 3.1 State Definition
where {G}, {I} and {C} are the global frame, local frame and camera frame, respectively. ${{}^{G}\textbf{p}_{I}}$ and ${{}^{G}\textbf{v}_{I}}$ are the position and velocity of IMU expressed in {G}, respectively. ${{}^{G}_{I}\textbf{q}}$ represents the rotation quaternion from {I} to {G} (in this paper, quaternion obeys Hamilton rules [29]). The vectors $\textbf{b}_{a}$ and $\textbf{b}_{g}$ individually represent the biases of the angular velocity and linear acceleration measured by the IMU device. And thus the corresponding EKF error-state of ${\textbf{x}_{I}}$ is defined as Eq. 2
@@ 3.2 Propagation and Augmentation
$\begin{matrix}{}^{G}_{I}\dot{\hat{\textbf{q}}}=\frac{1}{2}{{}^{G}_{I}{\hat{\textbf{q}}}}\otimes\Omega(\hat{\boldsymbol{\omega}}),&\Omega{(\hat{\boldsymbol{\omega}})}=\begin{pmatrix}0&-{\hat{\boldsymbol{\omega}}}^{\mathsf{T}}\\ \hat{\boldsymbol{\omega}}&-[\hat{\boldsymbol{\omega}}]_{\times}\end{pmatrix}\end{matrix}$ (5)
@@
where $\hat{\boldsymbol{\omega}}=\boldsymbol{\omega}_{m}-\hat{\textbf{b}}_{g}$ and $\hat{\textbf{a}}=\textbf{a}_{m}-\hat{\textbf{b}}_{a}$ are IMU measurements with biases discarded. where $[\hat{\boldsymbol{\omega}}]_{\times}$ is skew symmetric matrix of $\hat{\boldsymbol{\omega}}$. Based on Eqs.(5) to (7), the linearized continuous dynamics for the error IMU state is defined as
@@ 3.3 Schur Complement-Based State Update
$\begin{matrix}\textbf{r}=\begin{bmatrix}\textbf{J}_{x}&\textbf{J}_{f}\end{bmatrix}\begin{bmatrix}{\tilde{\textbf{X}}}\\ {{}^{G}{\tilde{\textbf{p}}_{f}}}\end{bmatrix}+\textbf{n}\end{matrix}$ (19)
where, r and $\begin{bmatrix}\textbf{J}_{x}&\textbf{J}_{f}\end{bmatrix}$ are respectively the stacked residual and stacked Jacobian. $\textbf{J}_{x}$ and $\textbf{J}_{f}$ are jacobian with respect to states and landmark positions, respectively. n is the stacked measurement noise, and the measurement covariance of n is $\textbf{R}=diag(u^{2},u^{2},\dotsb,u^{2})$, where $u$ is the element of standard deviation of n.
Unlike [30, 10, 7], in this paper, the residual model Eq. 19 is projected into the jacobian space $\begin{bmatrix}\textbf{J}_{x}&\textbf{J}_{f}\end{bmatrix}^{\mathsf{T}}$for formulating equivalent residual equations, which consist of gradient and hessian and observation covariance shown in Eqs. (20) and (21) below. It is worth highlighting that this strategy is an alternative to QR decomposition strategy [22] for speeding-up in any EKF systems with high-dimensional measurements.
@@ 4.1 Accuracy
In Tab. 1, our method obtains the lowest average RMSE in filter-based methods reported on the dataset so far, as well as outperforms the majority of optimization-based methods. Besides, our approach obtains the similar accuracy with wellknown optimization-based method BASALT and slightly lower accuracy than the recent competitor DM-VIO. Besides, the well-known VINS algorithms, VINS-Fusion [24] and SMSCKF [30], are not included in Tab. 1, since VINS-mono and OpenVINS surpass VINS-Fusion and SMSCKF in terms of accuracy, respectively [25, 10]. The re-evaluation experiment in Tab. 2 is within expectation absolutely. It is worth highlighting that, although degrading in accuracy slightly compared with the two optimization-based competitors, our method achieves obviously lower computational complexity than both of them with details in the next subsection.
@@ 4.2 Efficiency
As demonstrated in Tab. 3, SchurVINS achieves almost the lowest processor usage compared with all the mentioned VINS algorithms. Especially, SVO2.0-wo requires similar cpu usage with SchurVINS, but it suffers from notable inaccuracy since it is almost pure Visual Odometry (VO). To thoroughly investigate the underlying reasons contributing to the efficiency advantages of SchurVINS, the experiment to meticulously analyze the overhead time of SchurVINS including the comparison with SVO2.0, the widely-recognized filter-based OpenVINS and SMSCKF is carried out in Tab. 4.